OpenStudy (anonymous):

x-3 3x + 4 _____ - ________ = 3 2 2 x-3 3x + 4 _____ - ________ = 3 2 2 @Mathematics

OpenStudy (anonymous):

Multiply everything by 2. x-3-3x-4=6 Group everything and solve for x.

OpenStudy (anonymous):

Times everything by two (the numerator and denominator). The answer is as follows: \[\left( 2x-6-6x+8 \right)\div \left( 4 \right)=3\] Simplify: \[\left( -4x+2 \right)\div \left( 4 \right)=3\] Bring the 4 over to the other side of the equals sign, turning it into a times: \[-4x+2=12\] Finally: \[-4x=12-2\] \[-4x=10\] \[x=10\div\left( -4 \right)\] \[x=-2.5\] Hope this helps!